OpenStudy (anonymous):

The product of the page numbers on two facing pages of a book is 182. Find the page numbers.

OpenStudy (anonymous):

If two pages face each other, then they differ by only one. (How can two pages differ by more than 1 if they face each other? They would have a page between them.) So let n be the number of the first page. Then n+1 must be the number of the second page. The product of these two numbers, n*(n+1) is 182, Thus, n^2+n = 182. You can solve this for n (the first page number) by using the quadratic formula.

OpenStudy (zarkon):

182=2*7*13 so the page numbers must be 13 and 2*7=14

OpenStudy (anonymous):

Thank you, I was confused I was trying to figure out what the pages were if you added them together, forgot product meant multiplication

OpenStudy (anonymous):

Nice method, Zarkon. I hadn't even considered factoring it.